OpenStudy (anonymous):

Stupid Question... whats the vertex of y=x^2? Is it 0,0?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

whats the vertex of -2(x+3)^2+1 @Data_LG2

OpenStudy (anonymous):

y=a(x-h)^2+k --> (h,k) are the coordinates of the vertex

OpenStudy (anonymous):

so 0,3

OpenStudy (anonymous):

oh! 3,1

OpenStudy (anonymous):

Correct?

OpenStudy (anonymous):

which graph would open down? 0,0 or 3,1?

OpenStudy (anonymous):

if the value of a is negative then it's open down, if a is positive it's open up

OpenStudy (anonymous):

y=x^2 y=-2(x+3)^2+1 which one opens down and which one is more narrow? So the narrow is y=x^2 and opens down is -2(x+3)^2+1

OpenStudy (anonymous):

-2(x+3)^2+1 is more narrow

OpenStudy (anonymous):

thanks Data

OpenStudy (anonymous):

XD lol you already figure it out all by yourself... k gtg